# From Dataset to Real-world: General 3D Object Detection via Generalized Cross-domain Few-shot Learning

### v1

@@ 4.1 Image-guided Multi-modal Fusion
Figure 3: Physical-aware box searching, where red boxes are GT boxes and blue ones are searched boxes. Regarding “Cyclist” (left) and “Car” (right), searched boxes have biased angles and centers, respectively, which are corrected by $L_{BVC}$ and $L_{FVD}$.
@@
Furthermore, we notice that due to central, unidirectional scanning, LiDAR-based object points present significant differences in point distribution regarding different structural complexities. For simple structural objects with flat surfaces, mostly including vehicles (e.g., cars, trucks, and buses), most points are on the smooth surfaces, which are front-viewed by LiDAR. For complex structural objects with irregular surfaces (e.g., pedestrians, bicycles, and sitting persons), points are more to shape the whole objects in the bird’s view. Motivated by this observation, we categorize general objects into two types by their structural complexity: simple structural (SS) objects and complex structural (CS) objects. For SS objects, we design the front-viewed distance (FVD) loss to make points closer to the front-viewed boundaries of boxes,
$L_{FVD}=\sum_{\begin{subarray}{c}c=1\\ c\in SS\end{subarray}}^{C_{cls}}\sum_{\textbf{P}^{local}_{i}\in c}(|\textbf{P}^{local}_{i}|-\textbf{FB}^{c}),$ (6)
@@
where $\textbf{FB}^{c}$ is LiDAR-front-viewed box boundaries determined by the box heading angle $\theta$ and box location $[x,y,z]$ (refer to Supplementary for details). For CS objects, we design the bird-viewed center (BVC) loss to align the centers of points and boxes.
$L_{BVC}=\sum_{\begin{subarray}{c}c=1\\ c\in CS\end{subarray}}^{C_{cls}}\sum_{\textbf{P}^{local}_{i}\in c}|f_{2D}(\textbf{P}^{local}_{i})|,$ (7)
where $f_{2D}$ is the function to obtain $[x,y]$ of $\textbf{P}^{local}_{obj}$. Applying $L_{BVC}$ and $L_{FVD}$ facilitates finding correct centers and heading angles for boxes, as shown in Figure 3.
In summary, the loss of box searching for $[x,y,z,\theta]$ is:
$L_{box}=L_{OD}+\lambda_{1}L_{FVD}+\lambda_{2}L_{BVC}$ (8)
Since the computational load of box searching is low (due to the limited number of object points) and does not involve in model parameter update, we use the Quasi-Newton ”BFGS” function of the Scipy tool to quickly calculate the set of location $[x,y,z]$ and heading angle $\theta$ for each $\textbf{P}^{local}_{i}$.
@@
4.2 Contrastive-Enhanced Learnable Prototype and Feature Refinement
Our image-guided fusion significantly improves RPN proposal accuracy. Yet, the domain gap, sparse point density, and data limitation still prevent the model from learning reliable object features through simple model fine-tuning. To tackle this, we design a novel contrastive-enhanced prototype learning strategy from the limited samples. Refining proposal features with such prototypes enhances the model’s ability to detect objects-of-interest. Different from [18] that uses contrastive learning to enhances static prototypes, our approach leverages few-shot-driven contrastive learning on learnable prototypes, making our learnable prototypes more discriminative and comprehensive.
@@ 4.2 Contrastive-Enhanced Learnable Prototype and Feature Refinement
Figure 4: Few-shot feature extraction and CL-enhanced prototype learning. (In few-shot feature extraction, 2D and 3D ground-truth labels replace GDINO and RPN outputs to obtain object features.)
@@
Contrastive-enhanced prototype learning. We build a learnable targe-specific feature bank $\textbf{F}^{pro}\in\mathbb{R}^{|C^{t}|\times C}$ for all object classes. These prototypes are optimized together with the model finetuning update. To accelerate their converge under limited data constraint, we introduce a contrastive loss directly on these learnable prototypes. To this end, as shown in Figure 4, we pool few-shot object features $\textbf{F}^{fs}$ according to their ground-truth annotation as the contrastive anchors. Then for each class $c\in C^{t}$, we construct positive pairs with the corresponding prototype $\textbf{F}^{pro}_{c}$ and its anchor $\textbf{F}^{fs}_{c}$. The rest prototypes in the feature bank, denoted by $\textbf{F}^{pro}_{s}$ for $c\neq s$, are negative samples of the anchor.
$L_{CL}=-\sum_{c\in C_{cls}}\log\frac{\exp(\text{Sim}(\textbf{F}^{fs}_{c},\textbf{F}^{pro}_{c}))}{\sum_{s\in C_{cls}}\exp(\text{Sim}(\textbf{F}^{fs}_{c},\textbf{F}^{pro}_{s}))},$ (9)
@@
where $\text{Sim}(\cdot)$ calculates the cosine similarity between two numerical features in the InfoNCE contrastive loss [26]. Since the anchors are directly obtained from examples in the target domain, our contrastive-enhanced features helps bridge the domain gap between source and target environments and speed up $\textbf{F}^{pro}$ acquiring semantic essences of various classes under limited training data.
@@
Feature refinement by prototypes. After obtaining the $\textbf{F}^{pro}$ along with model finetuning process, we freeze the prototypes and use them to refine the proposal features $\textbf{F}^{prp}$ from the query data. In the multi-head cross-attention, we take $\textbf{F}^{pro}$ to form the key and value and $\textbf{F}^{prp}$ as the query.
$\hat{\textbf{F}}^{prp}=\text{Softmax}(\frac{{\textbf{F}}^{prp}\textbf{W}_{Q}({\textbf{F}}^{pro}\textbf{W}_{K})^{T}}{\sqrt{d}}){\textbf{F}}^{pro}\textbf{W}_{V},$ (10)
@@
where $d$ is the dimension of $\textbf{F}^{prp}$, and $\textbf{W}_{Q},\textbf{W}_{K},\textbf{W}_{V}$ are trainable transformations of query, key, and value. Finally,
$\tilde{\textbf{F}}^{prp}=\hat{\textbf{F}}^{prp}+\textbf{F}^{prp}.$ (11)
@@ 4.3 Model Optimization and Inference
The 3D detection model parameter update and our prototype learning are collectively conducted together. The overall loss $L$ to optimize the whole system is:
$L=L_{rpn}+L_{det}+{\lambda}L_{CL}$ (12)
where $L_{rpn}$ and $L_{det}$ are the standard detection loss of RPN and detection head, and $\lambda$ is a weight hyper-parameter. To further enable the model to efficiently adapt to a new domain under the limited data setting, we adopt a MAML-based [8] training scheme. Briefly, during meta-training, we leverage the source data to set up the $K$-shot meta-task.This meta-training facilitates to find a set of model parameters and $\textbf{F}^{pro}$ for the quick model adaptation in the unseen target domain. Please refer to the supplementary for more details.
During model deployment, the query point cloud and corresponding 2D image undergo the proposed image-guided fusion to enhance voxel representation and proposal boxes. After ROI pooling, the extracted features are further refined with class prototypes to improve discrimination.
@@ 5 Experimentation
Table 2: Performance in mAP(%) for NuScenes $\rightarrow 5$shot-KITTI, Waymo $\rightarrow 5$shot-KITTI, and KITTI $\rightarrow 5$shot-A2D2. The bold values represent the best performance. Subscript values in parentheses are standard deviations. Please refer to the supplementary for specific detection performance across various categories.
Methods Venues NuScenes→FS-KITTI Waymo→FS-KITTI KITTI→FS-A2D2
common novel overall common novel overall common novel overall
VoxelRCNN Target-FT - 12.77(1.85) 5.48(1.16) 8.61(1.45) 23.06(1.64) 12.47(1.92) 17.01(1.8) 5.09(1.1) 0.7(0.17) 2.9(0.64)
Proto-Vote [53] NIPS’22 7.56(2.39) 5.74(1.54) 6.52(1.9) 17.36(2.72) 12.08(1.78) 14.34(2.19) 3.61(0.81) 1.86(0.48) 2.74(0.65)
PVAE-Vote [37] NIPS’24 8.01(2.77) 6.38(2.22) 7.08(2.46) 18.19(2.94) 12.79(2.2) 15.1(2.52) 3.43(0.85) 1.97(0.54) 2.7(0.7)
CP-Vote [18] PRCV’24 10.69(2.28) 7.84(1.87) 9.06(2.04) 17.66(2.4) 12.17(1.86) 14.52(2.09) 4.28(0.92) 2.72(0.88) 3.5(0.9)
GFS-Det [19] arXiv’23 12.83(2.36) 1.18(0.38) 6.17(1.23) 22.74(2.78) 1.26(0.38) 10.47(1.41) 4.39(0.57) 0.22(0.05) 2.3(0.31)
Ours - 15.99(1.62) 11.72(1.4) 13.55(1.5) 25.4(2.02) 17.75(1.66) 21.03(1.82) 7.78(0.68) 5.22(0.61) 6.5(0.64)
PV-RCNN++ Target-FT - 15.28(2.03) 6.39(1.71) 10.2(1.85) 24.35(1.4) 13.79(1.73) 18.31(1.59) 3.64(1) 0.26(0.09) 1.95(0.55)
Proto-Vote [53] NIPS’22 6.97(1.43) 7.05(2.02) 7.02(1.76) 10.02(2.16) 12.27(1.49) 11.31(1.78) 3.92(0.81) 1.48(0.5) 2.7(0.65)
PVAE-Vote [37] NIPS’24 7.43(1.72) 7.53(2.16) 7.49(1.97) 11.36(2.4) 12.84(1.93) 12.21(2.13) 3.97(0.96) 1.34(0.54) 2.66(0.75)
CP-Vote [18] PRCV’24 8.73(1.97) 7.17(1.8) 7.84(1.87) 14.18(2.79) 13.32(2.74) 13.69(2.76) 4.2(1.1) 2.14(0.63) 3.17(0.87)
GFS-Det [19] arXiv’23 17.37(2.08) 1.16(0.14) 8.1(0.97) 24.79(1.79) 1.33(0.23) 9.67(0.9) 3.46(0.56) 0.14(0.04) 1.8(0.3)
Ours - 18.06(1.46) 11.11(1.03) 14.09(1.21) 27.35(1.55) 16.02(1.93) 20.88(1.77) 9.75(1.09) 5.76(0.8) 7.76(0.94)
5.1 Experimental Settings [^1] More details on experimental settings (e.g. benchmark setup, implementation details, and comparison methods) are provided in our supplementary.
@@
Implementation details. Voxel-absed VoxelRCNN [5] and point-voxel-based PVRCNN++ [33] are taken as the base detection models in experiments and our solution is implemented based on OpenPCDet [38] and RoboFusion [35]. Data augmentations during model pre-training include random world flipping, random world scaling, and random world rotation. In model fine-tuning, we additionally involve ground-truth object sampling to ensure the existence of all classes in each iteration. For box searching, we define CS classes [Pedestrian, Person_sitting, Cyclist, Bicycle, Utility_vehicle] and SS classes [Car, Truck, Van, Tram, Bus]. The “Adam-Onecycle” optimizer [38] with a learning rate of 0.01 is adopted. Epochs of pre-training are 30 for NuScenes and Waymo, and 80 for KITTI, the epoch of fine-tuning is 100 for both tasks. All experiments are conducted on a machine with dual-GeForce RTX-3090.
Compared Methods. As no prior work has specifically tackled the GCFS task for outdoor 3D object detection, we use a simple fine-tuning approach on target-domain data (denoted as Target-FT) as the baseline. To benchmark our method, we compare against SOTA 3D-FSL methods, including Proto-Vote [53], PVAE-Vote [37], and CP-Vote [18], as well as the 3D-GFSL method GFS-Det [19]. Note that current outdoor OVD methods (i.e., Unsup3D [24], FnP [6], and OpenSight [51]) and 3D-DA methods (i.e., SN [40] and ST3D++ [48, 47]) are not directly applicable to our GCFS benchmark, as they rely on extensive data, additional temporal information, or unsupervised learning without annotation-based supervision. To further assess the generalizability and potential of our approach, we extend our ablation study to a more complex unsupervised few-shot learning setting, where these 3D-OVD and 3D-DA methods can be evaluated under conditions more aligned with their original assumptions.
5.2 Experimental Results on GCFS Benchmark
For the $5$-shot GCFS tasks, our proposed method outperforms all compared approaches, as shown in Table 2. It achieves higher mAP for both common and novel categories [^3] The specific detection performance across various categories on the three benchmark settings is reported in the supplementary., demonstrating strong knowledge transferability from the source domain while effectively adapting to novel classes with limited supervision. Furthermore, our method exhibits robust adaptability across both high-density (64-beam) KITTI and low-density (16-beam) A2D2, ensuring robust few-shot detection even in challenging low-density LiDAR scenarios [17]. In contrast, existing 3D-FSL and 3D-GFSL methods struggle in GCFS tasks, primarily due to their limited ability to bridge domain gaps. For the 3D-FSL methods, their reliance on dense, close-range point clouds further degrades performance when detecting sparse, distant objects under varying lighting and weather conditions. GFS-Det particularly struggles with novel categories due to its simplistic incremental learning strategy, which prevents effective knowledge transfer from common classes to novel ones. In addition, we observe that all methods exhibit lower performance in the KITTI (64-beam) → FS-A2D2 (16-beam) setting. The significant reduction in LiDAR resolution introduces a larger domain gap, making domain adaptation more challenging compared to other settings [17].
Limitation on hard classes. Although our method outperforms other methods (see supplementary for details), it shows limited improvement on certain hard classes, (i.e., “Person_sitting”) due to the diverse and misleading structures of these samples, making feature learning more challenging. In the future, we plan to enhance our method to better handle these hard classes.
5.3 Ablation Studies
Table 3: Component ablations in mAP(%). (Image-Fusion is our proposed image-guided multi-modal fusion and CL-Proto is our proposed contrastive-learning-enhanced prototype learning.)
@@ 5.3 Ablation Studies
Table 4: Performance in mAP(%) of adopting prototype learning with or without contrastive learning (CL).

### v2

@@ 3.1 Image-guided Multi-modal Fusion (IMMF)
Figure 3: Physical-aware box searching. Red boxes are GT boxes, and blue ones are searched boxes. Regarding “Cyclist” (left) and “Car” (right), angle and center biases on searched boxes are corrected by $L_{\text{BVC}}$ and $L_{\text{FVD}}$.
Physical-aware 3D box searching from 2D masks. While VLMs provide rich semantics, transferring these 2D cues into the 3D space is inherently noisy in sparse LiDAR settings. Calibration inaccuracies and vision misalignment can lead to imprecise 2D-to-3D mappings. To ensure 2D semantic cues are projected to geometrically plausible 3D box proposals, we introduce a physically-aware box search strategy that filters and aligns proposals based on spatial consistency.
Specifically, to estimate fine-grained box locations from the 2D object masks $\textbf{M}^{\text{2D}}$, we first project the raw point cloud P into the image and identify points within masks by $\textbf{P}^{\text{pts}}=f_{\text{proj}}(\textbf{M}^{\text{2D}},\textbf{P})^{\top}\textbf{P}$, where $\textbf{P}^{\text{pts}}$ denotes the points of all object masks. For the $i^{th}$ object, we extract its points $\textbf{P}^{\text{pts}}_{i}\in\textbf{P}^{\text{pts}}$ and use the mean and $2\timesstandard$ deviation of point coordinates as the center and boundary of the valid range to eliminate background points. For each class $c\in\mathbb{C}^{t}$, we pre-define an anchor box with the size $h^{c},w^{c},l^{c}]$ via the mean size of target few-shot objects. The goal of box searching is to find the optimal center $[x,y,z]$ and heading angle $\theta$ of the anchor box for each object. Specifically, for $i$-th object, $x,y,z,\theta]$ defines a rotation transformation T (see the supplementary for details), and centered coordinates $\textbf{P}^{\text{local}}_{i}$ are obtained by $\textbf{P}^{\text{local}}_{i}=\textbf{T}\textbf{P}^{\text{pts}}_{i}$. We first design an outside distance loss $L_{\text{OD}}$ to constrain $\textbf{P}^{\text{local}}_{i}$ in the box,
$L_{OD}=\sum_{\textbf{p}\in\textbf{P}^{\text{local}}_{i}}\textrm{min}(\textrm{abs}(\textbf{p})-\textbf{BD}^{c},\textbf{0}).$ (3)
Here, $\textbf{BD}^{c}=[h^{c}/2,w^{c}/2,l^{c}/2]$ denotes the local box boundary for class $c$.
@@
Furthermore, we notice that, due to central unidirectional scanning, LiDAR-scanned object points present significant differences in point distribution regarding different structural complexities. For simple structural objects with flat surfaces, like vehicles (e.g., cars and buses), most points are on smooth surfaces and front-viewed by LiDAR. For complex structural objects with irregular surfaces (e.g., pedestrians and bicycles), points are more to shape the whole objects in the bird’s eye view. Motivated by this observation, we categorize general objects into two types: simple structural (SS) objects and complex structural (CS) ones. For SS objects, we design the front-viewed distance (FVD) loss to make points closer to the front-viewed boundaries of the box,
[…]
@@
where the LiDAR front-viewed box boundaries $\textbf{FB}^{c}$ is defined by $x,y,z,\theta]$ (see the supplementary for details). For CS objects, we design the bird-viewed center (BVC) loss to align the centers of points and boxes.
[…]
where $f_{\text{2D}}(\cdot)$ simply obtains $[x,y]$ of points. Applying $L_{\text{BVC}}$ and $L_{\text{FVD}}$ facilitates the discovery of the correct centers and heading angles for boxes, as shown in Fig. 3. In summary, the box-searching loss for optimizing $x,y,z,\theta]$ is:
$L_{\text{box}}=L_{\text{OD}}+\lambda_{1}L_{\text{FVD}}+\lambda_{2}L_{\text{BVC}}.$ (6)
Since the computational load of box searching is low (due to sparse object points), we use the Quasi-Newton BFGS optimization (Head and Zerner 1985) to efficiently optimize $x,y,z,\theta]$ for each object. In essence, our physically-aware box search acts as a semantic gatekeeper-ensuring that 2D-to-3D knowledge transfer remains spatially coherent.
@@
3.2 Class-specific Contrastive-Enhanced Learnable Prototype and Feature Refinement
While our IMMF module improves proposal accuracy, domain shifts and limited annotations still hinder reliable feature learning via simple fine-tuning. To overcome this, we propose a contrastive prototype learning strategy that builds robust, class-specific semantic anchors from limited examples and enhances them using contrastive learning to increase generalization and inter-class separability. Unlike the work (Li et al. 2024), which uses contrastive learning to enhance static prototypes, our approach uses few-shot-driven contrastive learning on learnable prototypes, making our prototypes more discriminative.
@@ 3.2 Class-specific Contrastive-Enhanced Learnable Prototype and Feature Refinement
Figure 4: Few-shot feature extraction and CL-enhanced prototype learning. In few-shot feature extraction, 2D and 3D ground-truth labels replace GDino and RPN outputs to extract object features.
@@
Class-specific contrastive prototype learning. We build a learnable target-specific feature bank $\textbf{F}^{\text{pro}}\in\mathbb{R}^{|\mathbb{C}^{t}|\times d}$ for all object classes, where $d$ is the dimension of features. These prototypes are optimized together with the model fine-tuning update. To accelerate convergence under limited data, we introduce a contrastive loss for the learnable prototypes. As shown in Fig. 4, we group the features of the few shots $\textbf{F}^{\text{fs}}$ according to their box annotation as contrastive anchors. Then for each class $c\in\mathbb{C}^{t}$, we construct positive pairs with the corresponding prototype $\textbf{F}^{\text{pro}}_{c}$ and its anchor $\textbf{F}^{\text{fs}}_{c}$. The remaining prototypes in the feature bank, denoted by $\textbf{F}^{\text{pro}}_{s}$, are negative samples of the anchor.
$L_{\text{CL}}=-\sum_{c\in\mathbb{C}^{t}}\log\frac{\exp(\text{Sim}(\textbf{F}^{\text{fs}}_{c},\textbf{F}^{\text{pro}}_{c})/\tau)}{\sum_{s\in\mathbb{C}^{t}}\exp(\text{Sim}(\textbf{F}^{\text{fs}}_{c},\textbf{F}^{\text{pro}}_{s})\tau)},$ (7)
@@
where $\text{Sim}(\cdot,\cdot)$ calculates the cosine similarity between two features in the InfoNCE loss (Oord et al. 2018) with a temperature $\tau$. Since the anchors are directly obtained from target-domain examples, our contrastive-enhanced features help bridge the domain gap between source and target environments and speed up $\textbf{F}^{\text{pro}}$ acquiring semantic essences of various classes under limited training data.
@@
Feature refinement by prototypes. After obtaining the $\textbf{F}^{\text{pro}}$ along with the model finetuning process, we use them to refine the proposal features $\textbf{F}^{\text{prp}}$ of the query input. In the multi-head cross-attention, we take $\textbf{F}^{\text{pro}}$ to form the key and value, and $\textbf{F}^{\text{prp}}$ as the query.
$\hat{\textbf{F}}^{\text{prp}}=\text{Softmax}(\frac{{\textbf{F}}^{\text{prp}}\textbf{W}_{\text{Q}}({\textbf{F}}^{\text{pro}}\textbf{W}_{\text{K}})^{\top}}{\sqrt{d}}){\textbf{F}}^{\text{pro}}\textbf{W}_{\text{V}},$ (8)
@@
where $\textbf{W}_{\text{Q}},\textbf{W}_{\text{K}},\textbf{W}_{\text{V}}$] is the trainable transformation of the query, key, and value. Finally,
$\tilde{\textbf{F}}^{\text{prp}}=\hat{\textbf{F}}^{\text{prp}}+\textbf{F}^{\text{prp}},$ (9)
@@ 3.3 Model Optimization and Inference
The model parameter update and our prototype learning are conducted together. The Overall loss to optimize them is:
$L=L_{\text{rpn}}+L_{\text{det}}+{\lambda}L_{\text{CL}},$ (10)
where $L_{\text{rpn}}$ and $L_{\text{det}}$ are the standard losses of RPN and detection head, and $\lambda$ is a weight hyper-parameter. To further enable the model’s adaptability to a new domain under limited data, we adopt an MAML-based (Finn et al. 2017) training scheme. Briefly, during meta-training, we leverage the source data to set up the $K$-shot meta-task. This meta-training facilitates finding a set of model parameters and $\textbf{F}^{\text{pro}}$ for the quick model adaptation in the unseen domain (see the supplementary for more details). During deployment, aligned point clouds and images undergo the proposed image-guided fusion to enhance semantic discovery in proposals. After ROI pooling, object features are further refined with class prototypes to improve discrimination.
Methods NuScenes→FS-KITTI Waymo→FS-KITTI KITTI→FS-A2D2 KITTI→FS-Argo2
common novel overall common novel overall common novel overall common novel overall
Source-only 14.24 - - 26.85 - - 3.81 - - 6.65 - -
Target-FT 12.77(1.9) 5.48(1.2) 8.61(1.5) 23.06(1.6) 12.47(1.9) 17.01(1.8) 5.09(1.1) 0.70(0.2) 2.90(0.6) 3.18(0.2) 0.62(0.1) 1.39(0.1)
Proto-Vote 7.56(2.4) 5.74(1.5) 6.52(1.9) 17.36(2.7) 12.08(1.8) 14.34(2.2) 3.61(0.8) 1.86(0.5) 2.74(0.7) 3.33(0.9) 0.90(0.4) 1.63(0.5)
PVAE-Vote 8.01(2.8) 6.38(2.2) 7.08(2.5) 18.19(2.9) 12.79(2.2) 15.10(2.5) 3.43(0.9) 1.97(0.5) 2.70(0.7) 3.10(1.0) 0.92(0.3) 1.58(0.5)
CP-Vote 10.69(2.3) 7.84(1.9) 9.06(2.0) 17.66(2.4) 12.17(1.9) 14.52(2.1) 4.28(0.9) 2.72(0.9) 3.50(0.9) 2.72(0.9) 0.93(0.4) 1.47(0.5)
GFS-Det 12.83(2.4) 1.18(0.4) 6.17(1.2) 22.74(2.8) 1.26(0.4) 10.47(1.4) 4.39(0.6) 0.22(0.1) 2.30(0.3) 6.11(0.1) 0.03(0.0) 1.86(0.0)
Ours 15.99(1.6) 11.72(1.4) 13.55(1.5) 25.40(2.0) 17.75(1.7) 21.03(1.8) 7.78(0.7) 5.22(0.6) 6.50(0.6) 6.71(0.2) 2.07(0.2) 3.46(0.2)
Full-Target 41.34 18.35 28.21 41.34 18.35 28.21 36.61 5.99 21.30 31.75 18.48 22.46
Table 2: Performance in mAP(%) of VoxelRCNN for NuScenes $\rightarrow 5$shot-KITTI, Waymo $\rightarrow 5$shot-KITTI, KITTI $\rightarrow 5$shot-A2D2, and KITTI $\rightarrow 5$shot-Argo2. The bold values represent the best performance except Full-Target. Subscript values in parentheses are standard deviations. Please refer to the supplementary for specifics across various categories.
@@ 4 Experimentation
4.1 Experimental Settings [^1] Details on the benchmark setup and implementation are provided in the supplementary linked in our GitHub repository.
@@
Implementation Details. We use VoxelRCNN (Deng et al. 2021) (voxel-based) and PV-RCNN++ (Shi et al. 2023) (point-voxel-based) as base detectors. Pre-training applies standard augmentations: random world flipping, scaling, and rotation. In fine-tuning, we additionally use ground-truth object sampling to ensure all target classes are present in each iteration. For box searching, we define SS classes [Car, Truck, Van, Tram, Bus, Construction_barrel, Large_vehicle, Sign] and CS classes [Pedestrian, Person_sitting, Cyclist, Bicycle, Utility_vehicle, Traffic_cone, Motorcycle]. The Adam-OneCycle optimizer (Team 2020; Song et al. 2024) is used with a 0.01 learning rate. All models are pre-trained for 30 epochs on NuScenes and Waymo, 80 epochs on KITTI, and fine-tuned for 100 epochs in FS-datasets. Batch sizes are 2 in pre-training and 1 in fine-tuning and testing.
Compared Methods. As no prior work has specifically tackled GCFS tasks for outdoor 3D object detection, we use a simple fine-tuning on few-shot target data (Target-FT) as the baseline. Source-only training and full target supervision (Source-only and Full-Target) serve as the performance with no and full adaptation. To benchmark our method, we compare against SOTA 3D-FSL methods, Proto-Vote (Zhao and Qi 2022), PVAE-Vote (Tang et al. 2024), and CP-Vote (Li et al. 2024), as well as the 3D-GFSL method GFS-Det (Liu et al. 2023). Note that current outdoor OVD methods (i.e., Unsup3D (Najibi et al. 2023), FnP (Etchegaray et al. 2024), and OpenSight (Zhang et al. 2025a)) and 3D-DA methods (i.e., SN (Wang et al. 2020b), ST3D++ (Yang et al. 2022, 2021), and DenResamp (Li et al. 2025)) are not directly applicable to our GCFS benchmark, as they rely on extensive unannotated data for unsupervised learning. To further assess the generalizability and potential of our approach, we extend our ablation study to a more complex unsupervised few-shot learning setting, where these 3D-OVD and 3D-DA methods can be evaluated under conditions more aligned with their original assumptions.
@@
Table 3: Component ablations in mAP(%). Image-Fusion is our proposed IMMF module and CL-Proto is our proposed contrastive-learning-enhanced prototype learning.
4.2 Experimental Results on GCFS Benchmark
[…]
Limitations and Future Work. Our method shows limited gains on certain hard classes (e.g. “Person_sitting”) due to ambiguous and diverse structures. In low-shift scenarios without semantic changes (Waymo $\rightarrow$ FS-KITTI), improvements on common classes are marginal, due to the interruption of novel classes. Future work will focus on hard class learning, adaptability in shiftless settings, and code optimization for computation speed-up.
4.3 Ablation Studies